OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 4) = 2x. 2 -2 -8 -4

OpenStudy (tester97):

x=-2 and this is history not math:)

OpenStudy (anonymous):

thanks :3 can you help with a few more?

OpenStudy (tester97):

ok sure

OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 10 0 6 All Real Numbers No Solution

OpenStudy (tester97):

no sol.

OpenStudy (anonymous):

Kyra observed that 85% of the students at a library liked reading mystery stories. If 30 students at the library did not like reading mystery stories, the number of students at the library who liked mystery stories was _____. (only put numeric values, no other symbols) Numerical Answers Expected!

OpenStudy (tester97):

170

OpenStudy (anonymous):

The sum of two consecutive odd integers is 244. What is the smaller integer? 119 125 123 121

OpenStudy (tester97):

121 and 123

OpenStudy (tester97):

smaller is 121

OpenStudy (anonymous):

In a certain county, the number of charter schools is 6 less than twice the number of alternative schools. We know that there are 52 charter schools in the county. How many alternative schools are in the county? 58 44 29 23

OpenStudy (tester97):

29

OpenStudy (anonymous):

thanks :3 The length of a rectangle is 24 units. Can the perimeter x of the rectangle be 60 units when its width y is 11 units? No, the rectangle cannot have x = 60 and y = 11 because x = 48 + 2y. No, the rectangle cannot have x = 60 and y = 11 because x = 24 + 2y. Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 48. Yes, the rectangle can have x = 60 and y = 11 because x + y is greater than 24.

OpenStudy (tester97):

a

OpenStudy (anonymous):

The table below shows two equations: Equation 1 |2x – 3| + 5 = 4 Equation 2 |5x + 3| − 10 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = 2, -3.2. The solutions to equation 1 are x = 1, 2 and equation 2 has no solution. The solutions to equation 1 are x = 1, 2 and equation 2 has solutions x = 2, -3.2.

OpenStudy (tester97):

b

OpenStudy (anonymous):

thx Below are the steps to solve an equation: Step 1: |x – 5| + 2 = 5 Step 2: |x – 5| = 5 – 2 Step 3: |x – 5| = 3 Which of the following is a correct next step to solve the equation? x + 5 = -3 -x – 5 = 3 x - 5 = -3 x + 5 = 3

OpenStudy (tester97):

c?

OpenStudy (anonymous):

thanks he table shows the solution to the equation |2x – 3| - 1 = 2: Step 1 |2x – 3| = 2 + 1 Step 2 |2x – 3| = 3 Step 3 2x – 3 = 3 or 2x + 3 = 3 Step 4 2x = 6 or 2x = 0 Step 5 x = 3 or x = 0 Which is the first incorrect step? Step 1 Step 3 Step 5 Solution is correct

OpenStudy (tester97):

3

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 1) less than or greater to 10? x 6 6 x A number line with a closed circle on 6 and shading to the left A number line with a closed circle on 6 and shading to the right

OpenStudy (tester97):

d x<_ 6

OpenStudy (anonymous):

answers D?

OpenStudy (tester97):

i meant 

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 3(2x - 1) greater than or equal to 4(2x - 3) - 3? x less than or greater to 6 6 greater than or equal to x A number line with a closed circle on 6 and shading to the left A number line with a closed circle on 6 and shading to the right

OpenStudy (tester97):

c

OpenStudy (tester97):

thats all i can help wit sorry

OpenStudy (anonymous):

just a few more? :D